Session: Feb 2011 (session 1)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Effelsberg (Uwe Bach at Wed Mar 09 15:01:58 2011) Success:

Yebes (Pablo de Vicente / Paco Colomer at Wed Mar 09 15:06:59 2011) Minor failures:
Yebes data recorded on diskpack OAN+0020/8000/1024 (VSN label missing at start of observations). Some RFI seen at times in BBC04 LSB.

Torun (Kaz Borkowski at Wed Mar 09 15:25:28 2011) Minor failures: RECO
Inserting second diskpack (DP) into bank B at a break between scans at 8:06 UT made Mark5A report a 'Record pointer not increasing'. We had to reboot it twice and remove the DP from bank B to resume the recording at 9:09 only. Similar problem occured at DP change (11:18), when we again lost about 10 minutes of recording.

Westerbork (Geert Kuper at Wed Mar 09 15:38:36 2011) Success:
No known problems

Medicina (A. Cattani at Wed Mar 09 15:40:29 2011) Success:

Shanghai (Bo Xia at Fri Mar 11 08:14:13 2011) Success:

Onsala (Michael Lindqvist at Sat Mar 12 04:55:26 2011) Minor failures:
Problems with the Mk5 around 00:20 UTC, see log for details.

Urumqi (Zhang hua at Mon Mar 14 02:32:06 2011) Success:

Jodrell_Bank (Alastair Gunn at Tue Mar 29 11:33:53 2011) Minor failures: WIND
Telescope was parked between about 09:38 and 12:53 due to high winds. No other known problems.